# Improving Small Drone Detection Through Multi-Scale Processing and Data Augmentation

### v1

@@ III-B Proposed Approach
To optimize detection performance, a confidence threshold of 0.375 was experimentally established. To mitigate missed detections, a post-processing stage was implemented. Specifically, this routine examines each frame’s detected objects and searches for matching objects within a temporal window spanning six frames both preceding and following the target frame. The matching criteria require objects to share the same classification label and have a minimum degree of overlap (IoU $\geq$ 0.1). Upon identifying a potential match, the bounding box of the interpolated object is computed by linearly interpolating the bounding box coordinates between the matched objects. To ensure reliability, objects near image boundaries are discarded to prevent extrapolating detections for objects that have left the frame. Furthermore, the interpolated object is accepted only if it does not significantly overlap with existing detections in the target frame, avoiding redundant predictions. The confidence score of the interpolated object is computed as the mean confidence of the matching objects, reduced by a factor of 2 to account for the uncertainty introduced by interpolation. As illustrated in Fig. 4 (right), this process helps bridge detection gaps, improving temporal consistency and overall accuracy, albeit with a minor increase in the number of false positives.
Another post-processing strategy we explored involved incorporating a classifier to refine the regions detected by YOLO11m, aiming to further reduce false positives. We experimented with various models, ranging from lightweight networks designed for specific tasks [38, 39] to well-established architectures such as MobileNetV3 [40]. To minimize the impact on inference time, we prioritized smaller models. However, regardless of the model used, performance declined. While this approach effectively filtered out many false positives, it also resulted in the loss of many true positives (in other words, several drones correctly detected by YOLO11m were erroneously discarded by the classifier).
We chose not to implement a maximum-drone-count heuristic, as utilized in [27], due to its limited real-world applicability. While such heuristics can be effective in competition environments with short, pre-defined videos, they struggle to adapt to real-time video streams’ dynamic and unpredictable nature.
@@ III-C Results
We evaluated our approach on the validation set, which consists of seven videos from the Drone-vs-Bird challenge dataset, covering diverse conditions as described in Section III-A (the challenge organizers maintained exclusive access to the test set annotations). Performance was assessed using the mean Average Precision (mAP) at an IoU threshold of 50% (mAP 50). This metric calculates the average precision across recall values, considering a detection successful if the IoU between the predicted and ground truth bounding boxes is at least 50%.
The results, presented in Table II, show the mAP 50 achieved for each video. We compare two approaches: our full proposed method, which employs a multi-scale input processing strategy analyzing both whole images and segmented components, and a simplified variant processing only whole images.
@@
Table II clearly shows that processing the input images both as a whole and in segmented components improves detection performance, particularly for smaller drones. However, the dji_phantom_4_hillside_cross video stands as an exception, where processing only whole images resulted in better outcomes. Upon further analysis, we determined that the reason for this was the video’s complex hillside background, characterized by distant rocks and vegetation that closely mimic small-scale drones. When examining cropped bounding boxes from these regions, we occasionally mistook them for drones as well. In this particular video, the drone is white, providing a strong contrast against the background, and is relatively large. As a result, dividing the input image into segments was not strictly necessary. While incorporating segmented components in the processing still improved Recall, it resulted in a disproportionately larger decrease in Precision.
Examples of successfully detected drones are shown in Fig. 6. It is important to observe that the proposed approach is capable of detecting drones across various backgrounds, weather conditions, scales, among other factors. Of particular interest is the left image in the third row of Fig. 6 (also featured on the first page of this work), where none of the birds were erroneously identified as drones.
@@
Figure 6: Examples of drones successfully detected by the proposed approach. We have intentionally refrained from providing zoomed-in views of distant, small drones to emphasize the difficulty of detecting them.
@@ IV Conclusions
In this article, we describe our approach for drone detection, which ranked among the top 3 teams in the 8th WOSDETC Drone-vs-Bird Detection Grand Challenge [12].
We utilized the medium-sized YOLOv11 model [15] for drone detection. To overcome the model’s struggles with small-scale drone detection, we implemented a multi-scale approach in which the input image is processed both as a whole and in segmented components. This strategy significantly boosted detection performance, especially for distant drones.
@@
We aim to improve detection performance through the following investigations: (i) leveraging a state-of-the-art tracker to enhance multi-frame predictions; (ii) developing a classifier that analyzes multi-frame image patches to accurately differentiate drones from similar objects, such as birds, thereby reducing false positives [26]; and (iii) investigating latest architectures, such as D-FINE [41] and RT-DETRv2 [42], while incorporating a wider variety of datasets, including the recently introduced SynDroneVision [43], into the training process.
@@ ACKNOWLEDGMENTS
This study was supported in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, and by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under grants # $315409$/$2023$-$1$ and # $312565$/$2023$-$2$. We gratefully acknowledge the Pontifícia Universidade Católica do Paraná and Fundação Araucária for their financial support enabling conference participation. We also thank NVIDIA Corporation for donating the Quadro RTX $8000$ GPU used in this research.

### v2

@@ III-B Proposed Approach
To optimize detection performance, a confidence threshold of 0.375 was experimentally established. To mitigate missed detections, a post-processing stage was implemented. Specifically, this routine examines each frame’s detected objects and searches for matching objects within a temporal window spanning six frames both preceding and following the target frame. The matching criteria require objects to share the same classification label and have a minimum degree of overlap (IoU $\geq$ 0.1). Upon identifying a potential match, the bounding box of the interpolated object is computed by linearly interpolating the bounding box coordinates between the matched objects. To ensure reliability, objects near image boundaries are discarded to prevent extrapolating detections for objects that have left the frame. Furthermore, the interpolated object is accepted only if it does not significantly overlap with existing detections in the target frame, avoiding redundant predictions. The confidence score of the interpolated object is computed as the mean confidence of the matching objects, reduced by a factor of 2 to account for the uncertainty introduced by interpolation. As illustrated in Fig.˜4 (right), this process helps bridge detection gaps, improving temporal consistency and overall accuracy, albeit with a minor increase in the number of false positives.
Another post-processing strategy we explored involved incorporating a classifier to refine the regions detected by YOLO11m, aiming to further reduce false positives. We experimented with various models, ranging from lightweight networks designed for specific tasks [laroca2021towards, laroca2022first] to well-established architectures such as MobileNetV3 [howard2019searching]. To minimize the impact on inference time, we prioritized smaller models. However, regardless of the model used, performance declined. While this approach effectively filtered out many false positives, it also resulted in the loss of many true positives (in other words, several drones correctly detected by YOLO11m were erroneously discarded by the classifier).
We chose not to implement a maximum-drone-count heuristic, as utilized in [mistry2023drone], due to its limited real-world applicability. While such heuristics can be effective in competition environments with short, pre-defined videos, they struggle to adapt to real-time video streams’ dynamic and unpredictable nature.
@@ III-C Results
We evaluated our approach on the validation set, which consists of seven videos from the Drone-vs-Bird challenge dataset, covering diverse conditions as described in Section˜III-A (the challenge organizers maintained exclusive access to the test set annotations). Performance was assessed using the mean Average Precision (mAP) at an IoU threshold of 50% (mAP 50). This metric calculates the average precision across recall values, considering a detection successful if the IoU between the predicted and ground truth bounding boxes is at least 50%.
The results, presented in Table˜II, show the mAP 50 achieved for each video. We compare two approaches: our full proposed method, which employs a multi-scale input processing strategy analyzing both whole images and segmented components, and a simplified variant processing only whole images.
@@
Table˜II clearly shows that processing the input images both as a whole and in segmented components improves detection performance, particularly for smaller drones. However, the dji_phantom_4_hillside_cross video stands as an exception, where processing only whole images resulted in better outcomes. Upon further analysis, we determined that the reason for this was the video’s complex hillside background, characterized by distant rocks and vegetation that closely mimic small-scale drones. When examining cropped bounding boxes from these regions, we occasionally mistook them for drones as well. In this particular video, the drone is white, providing a strong contrast against the background, and is relatively large. As a result, dividing the input image into segments was not strictly necessary. While incorporating segmented components in the processing still improved Recall, it resulted in a disproportionately larger decrease in Precision.
Examples of successfully detected drones are shown in Fig.˜6. It is important to observe that the proposed approach is capable of detecting drones across various backgrounds, weather conditions, scales, among other factors. Of particular interest is the left image in the third row of Fig.˜6 (also featured on the first page of this work), where none of the birds were erroneously identified as drones.
@@
Figure 6: Examples of drones successfully detected by the proposed approach. In this figure, we have intentionally refrained from providing zoomed-in views of distant, small drones to highlight the difficulty of detecting them.
@@ IV Conclusions
In this article, we describe our approach for drone detection, which achieved first place in the 8th WOSDETC Drone-vs-Bird Detection Grand Challenge [challenge].
We utilized the medium-sized YOLOv11 model [yolov11] for drone detection. To overcome the model’s struggles with small-scale drone detection, we implemented a multi-scale approach in which the input image is processed both as a whole and in segmented components. This strategy significantly boosted detection performance, especially for distant drones.
@@
We aim to improve detection performance through the following investigations: (i) leveraging a state-of-the-art tracker to enhance multi-frame predictions; (ii) developing a classifier that analyzes multi-frame image patches to accurately differentiate drones from similar objects, such as birds, thereby reducing false positives [akyon2022sequence]; and (iii) investigating latest architectures, such as D-FINE [peng2024dfine] and RT-DETRv2 [lv2024rtdetrv2], while incorporating a wider variety of datasets, including the recently introduced SynDroneVision [lenhard2025syndronevision], into the training process.
@@ ACKNOWLEDGMENTS
This study was supported in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, and by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under grants # 315409315409/20232023-11 and # 312565312565/20232023-22. We gratefully acknowledge the Pontifícia Universidade Católica do Paraná and Fundação Araucária for their financial support enabling conference participation. We also thank NVIDIA Corporation for donating the Quadro RTX $8000$ GPU used in this research.